a birthday balloon is filled to a volume of 1.5L of helium gas in an air conditioned room ay 293k. the balloon is taken outdoors on an warm day where the volume expands to 1.55L. Assuming the pressure temperature outside in Celsius.


1
Expert's answer
2021-10-12T02:05:03-0400

V1/T1= V2/T2

1.5/293= 1.55/T2

0.00512T2= 1.55

T2 = 303.92k

= 30.92°C